OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 + 18x + 71 y = (x - 9)^2 - 10 y = (x + 9)^2 - 10 y = (x + 9)^2 + 10 y = (x - 9)^2 + 10

hartnn (hartnn):

so its basically completing the square, have heard of it?

OpenStudy (anonymous):

I've heard of it , yes

hartnn (hartnn):

ok, in this method, you add and subtract this term, "[(1/2)*co-efficient of x]^2" so, whats the co-efficient of 'x' here?

OpenStudy (anonymous):

1

hartnn (hartnn):

umm...no, co-efficient of 'x' is the number multiplied with 'x' (not x^2)

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

18?

hartnn (hartnn):

yes! 1/2*18 =...?

hartnn (hartnn):

and its square = ...?

OpenStudy (anonymous):

wait what?

hartnn (hartnn):

ok, step by step :P 1/2 *18 = ...?

OpenStudy (anonymous):

9

hartnn (hartnn):

yes, 2nd last step : 9^2= ... ?

OpenStudy (anonymous):

81

hartnn (hartnn):

last step: retriceand subtract 81 to y = x^2 + 18x + 71 y = (x^2 + 18x +81) -81+71 what can you tell about (x^2 + 18x +81) ?

hartnn (hartnn):

i meant "add and subtract"

OpenStudy (anonymous):

I don't know :( i'm lost

hartnn (hartnn):

did you get this? v

hartnn (hartnn):

y = (x^2 + 18x +81) -81+71

OpenStudy (anonymous):

yes .

hartnn (hartnn):

ok, so you know \(a^2+2ab+b^2=(a+b)^2\) ?

OpenStudy (anonymous):

kinda , yeah

hartnn (hartnn):

can i say \((x^2 + 18x +81) = (x+9)^2\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, which choice will you choose now ?

OpenStudy (anonymous):

c?

hartnn (hartnn):

\(y = (x^2 + 18x +81) -81+71=(x+9)^2-10\) not b?

OpenStudy (anonymous):

yeah b , sorry I got them mixed up

hartnn (hartnn):

no porbelm :)